3. 1
------- + -------
X - 2. 2 - x

Answer by stanbon(75887)

3. 1
------- + -------
X - 2. 2 - x
----------------------------
Note: -(x-2) = 2-x
==========================
Rearrange:
[3/(x-2) -[1/(x-2)]
----
= (3-1)/(x-2)
---
= 2/(x-2)
==============
Cheers,
Stan H.